ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46 Yost, David, Chulalongkorn University, Thailand Twisted sums of classical Banach spaces Fix two Banach spaces S and Q. Given a Banach space X containing S for which X/S is isomorphic to Q, need X be isomorphic to the direct sum of S and Q? We consider this problem for the case when S and Q are classical function spaces. We are particularly interested in the case when S = c_{0} and Q is not separable.